# Existence and qualitative behavior of solutions of abstract differential-algebraic equations

### v1

@@ 3 Nonlocal (global) implicit function theorems
The proof of the existence of a unique solution of equation (3.1) with respect to $y$ in a certain neighbourhood of a point $(t^{\prime},x^{\prime},y^{\prime})$ satisfying the equation is similar to the proof of the implicit function theorem without the requirement of differentiability [31, p. 420]. The Lipschitz continuity of the implicit function with respect to $x$ follows from the local Lipschitz continuity of $F$ with respect to $(x,y)$ and the contractivity of the mapping $\widehat{F}$ defined in (3.3) below. Note that $U_{r_{1}}(t^{\prime})$ can be a semi-open interval, as indicated in the introduction. In general, the idea of the proof is the same as for [10, Lemma 3.1]. However, since in [10] nonregular (singular) semilinear DAEs in finite-dimensional spaces were considered, then the form of components of the phase variable $x$, as well as direct decompositions of spaces and the associated projectors, is different from that of the present paper. The proof of [10, Lemma 3.1] is quite cumbersome due to the larger number of arguments of an implicit function and some features of the problem, so it will be easier to follow the proof of the lemma given above. Thus, for clarity, we provide the complete proof of the present lemma.
@@
It follows from the foregoing that, by the Banach’s fixed point theorem and the theorem on continuous dependence of a fixed point on a parameter (see, e.g., [28, Theorems 46, $46_{2}$]), the mapping $\widehat{F}(t,x,y)$ as a function of $y$, depending on the parameter $(t,x)$, has a unique fixed point $\upsilon_{(t,x)}=\upsilon(t,x)$ (i.e., $\widehat{F}(t,x,\upsilon(t,x))=\upsilon(t,x)$) in $\overline{U_{\tilde{\varepsilon}}}(y^{\prime})$ for each $(t,x)\in\overline{U_{\delta_{*}}}(t^{\prime},x^{\prime})$, which satisfies the equality $\upsilon(t^{\prime},x^{\prime})=y^{\prime}$ and depends continuously on $(t,x)$. The continuity of the function $\upsilon\colon\overline{U_{\delta_{*}}}(t^{\prime},x^{\prime})\to\overline{U_{\tilde{\varepsilon}}}(y^{\prime})$ is proved in the same way as in [28, Theorem $46_{2}$].
@@
Choose an arbitrary fixed point $(t^{\prime},x^{\prime})\in{\mathcal{T}}\times M_{X}$. Then, due to conditions 1 and 2, there exists a unique $y^{\prime}\in M_{Y}$} such that $(t^{\prime},x^{\prime},y^{\prime})$ satisfies (3.1) and the operator $\partial_{y}F(t^{\prime},x^{\prime},y^{\prime})$ has the bounded inverse $\partial_{y}F(t^{\prime},x^{\prime},y^{\prime})]^{-1}\in\mathrm{L}(Z,Y)$. Using the theorems on the existence and differentiability of an implicit function [28, Theorems 25, 28], we obtain that there exist open neighborhoods $U_{\delta_{1}}(t^{\prime})$, $U_{\delta_{2}}(x^{\prime})$ and $U_{\varepsilon}(y^{\prime})$ and a unique function $\upsilon_{t^{\prime},x^{\prime}}\in C\big(U_{\delta_{1}}(t^{\prime})\times U_{\delta_{2}}(x^{\prime}),U_{\varepsilon}(y^{\prime})\big)$ such that $\upsilon_{t^{\prime},x^{\prime}}(t^{\prime},x^{\prime})=y^{\prime}$, $\upsilon_{t^{\prime},x^{\prime}}(t,x)$ is a solution of equation (3.1) with respect to $y$, and it is continuously differentiable with respect to $x$ on $U_{\delta_{1}}(t^{\prime})\times U_{\delta_{2}}(x^{\prime})$. Recall that $U_{\delta_{1}}(t^{\prime})$ can be a semi-open interval in $\mathcal{T}$}. Generally, the implicit function theorems [28] assume that the set of variables is open, and therefore, to prove the existence of an implicit function when the interval $U_{\delta_{1}}(t^{\prime})$ is semi-open, the fixed point theorems [28, Theorems 46, $46_{2}$] as well as the proofs of the implicit function theorems [28, Theorems 25, 28] (which in turn use the same fixed point theorems) are used.
@@
Let us show this. Since $F(t,x,y)$ has a continuous partial derivatives with respect to $(x,y)$ on ${\mathcal{T}}\times D_{X}\times D_{Y}$, then it is locally Lipschitz continuous with respect to $(x,y)$. Choose an arbitrary fixed point $(t^{\prime},x^{\prime},y^{\prime})\in{\mathcal{T}}\times M_{X}\times M_{Y}$ at which equation (3.1) holds, and take the operator $W_{t^{\prime},x^{\prime},y^{\prime}}=\partial_{y}F(t^{\prime},x^{\prime},y^{\prime})$ as the operator $W_{t^{\prime},x^{\prime},y^{\prime}}$ appearing in condition 2 of Theorem 3.1. Then, due to condition 2 of Theorem 3.2, there exist open neighborhoods $U_{\delta}(t^{\prime},x^{\prime})\subset{\mathcal{T}}\times D_{X}$ and $U_{\varepsilon}(y^{\prime})\subset D_{Y}$ such that for each $(t,x)\in\overline{U_{\delta}(t^{\prime},x^{\prime})}$, $y^{1},\,y^{2}\in\overline{U_{\varepsilon}(y^{\prime})}$ the following holds: $\big\|F(t,x,y^{1})-F(t,x,y^{2})-W_{t^{\prime},x^{\prime},y^{\prime}}\big[y^{1}-y^{2}\big]\big\|\leq\int\limits_{0}^{1}\big\|\partial_{y}F\big(t,x,y^{2}+\theta(y^{1}-y^{2})\big)-W_{t^{\prime},x^{\prime},y^{\prime}}\big\|d\theta\,\|y^{1}-y^{2}\|\leq k(\delta,\varepsilon)\|y^{1}-y^{2}\|$, where $k(\delta,\varepsilon)=\sup\limits_{(t,x)\in\overline{U_{\delta}(t^{\prime},x^{\prime})},\,y^{\prime\prime}\in\overline{U_{\varepsilon}(y^{\prime})}}\big\|\partial_{y}F(t,x,y^{\prime\prime})-W_{t^{\prime},x^{\prime},y^{\prime}}\big\|\to 0$ as $\delta,\varepsilon\to 0$ since $\partial_{y}F$ is continuous at $(t^{\prime},x^{\prime},y^{\prime})$ (cf. [31, p. 422, the proof of the theorem]). Hence, taking into account that $W_{t^{\prime},x^{\prime},y^{\prime}}$ has a bounded inverse, condition 2 of Theorem 3.1 holds.
@@
The proof is similar to the proof of Theorem 3.2, except that the theorem on higher derivatives of an implicit function [28, Theorem 31] is also used here. ∎
Another nonlocal (global) implicit function theorems have been presented, e.g., in [1, 23, 3] (see also references therein). In [1], a nonlocal implicit function theorem for a closed mapping with a parameter from one Asplund space to another has been obtained in terms of the regular coderivative of the mapping. In [23], nonlocal implicit function theorems for mappings in Banach spaces has been obtained by using the concept of ‘‘continuation property’’ and local mapping relations. For mappings in real finite-dimensional spaces, a nonlocal implicit function theorem is presented in [3] with the use of the maximal rank condition.
@@ 4.1 First approach
In this subsection, we consider the case when the component $x_{20}=P_{20}x$ can be defined as an implicit function from equation (4.2) given below (i.e., when the projection $Q_{2*}f(t,x)$ depends on $x_{20}$). In the case when the pencil $P(\lambda)$ has index 1, this always holds, since $D_{20}=D_{2}$, $x_{20}=x_{2}=P_{2}x$, $x_{2\Sigma}=0$, $Q_{2*}=Q_{2}$ ($P_{2\Sigma}=0$, $Q_{2\Sigma}=0$) and $B_{2}=(Q_{2}B)\big|_{D_{2}}$ has the inverse $B_{2}^{-1}\in\mathrm{L}(Y_{2},D_{2})$.
@@
$F_{2*}(t,x_{1},x_{2\Sigma},x_{20})=0,\quad F_{2*}(t,x_{1},x_{2\Sigma},x_{20}):=Q_{2*}f(t,x)-Q_{2*}B(x_{20}+x_{2\Sigma}),$ (4.4)
where the operator $\widetilde{A}^{-1}\in\mathrm{L}(Y,D)$ is the inverse of
@@
and $A^{(-1)}$ is the semi-inverse operator to $A$, which can be obtained by $A^{(-1)}=\widetilde{A}^{-1}(Q_{1}+Q_{2\Sigma})$} (or by using the relations mentioned below). The existence and continuity of $\widetilde{A}^{-1}$ is proved in a similar way as the Schmidt lemma (see [31, p. 232] or [32]). It is readily verified that the operator $\widetilde{A}^{-1}=A^{(-1)}+\sum\limits_{i=1}^{n}\langle\cdot,q_{i}^{m_{i}}\rangle\varphi_{i}^{1}$ is the inverse of $\widetilde{A}$. Notice that $Y_{1}=\widetilde{A}D_{1}=AD_{1}$,  $Y_{2\Sigma}=\widetilde{A}D_{2\Sigma}=AD_{2\Sigma}$,  $\widetilde{A}^{-1}Ax=(P_{1}+P_{2\Sigma})x$, $x\in D$, and $A\widetilde{A}^{-1}y=(Q_{1}+Q_{2\Sigma})y$, $y\in Y$. In addition, $Q_{1}\widetilde{A}x=Q_{1}Ax=AP_{1}x=\widetilde{A}P_{1}x$ and $Q_{2\Sigma}\widetilde{A}x=Q_{2\Sigma}Ax=AP_{2\Sigma}=\widetilde{A}P_{2\Sigma}x$, $x\in D$.
Generally, $A^{(-1)}$ is defined by the relations $^{(-1)}A=P_{1}+P_{2\Sigma}, AA$$A^{(-1)}=Q_{1}+Q_{2\Sigma}$, $A^{(-1)}=(P_{1}+P_{2\Sigma})A^{(-1)}$ (or $A^{(-1)}=A^{(-1)}(Q_{1}+Q_{2\Sigma})$ ).
@@
where $\widetilde{A}=A+BP_{2}=Q_{1}A+Q_{2}B$ coincide with the operator $G=Q_{1}A+Q_{2}B$ from [33, Subsection 3.3] which has the inverse $G^{-1}=\widetilde{A}^{-1}\in\mathrm{L}(Y,D)$.
@@
The ‘‘algebraic part’’ (4.2) of the ADAE generates the manifold
@@
Since $\widetilde{A}^{-1}\in\mathrm{L}(Y,D)$, $Q_{1}+Q_{2\Sigma}\in\mathrm{L}(Y)$, $(Q_{1}+Q_{2\Sigma})B\widetilde{A}^{-1}\in\mathrm{L}(Y)$ and $(Q_{1}+Q_{2\Sigma})BP_{20}\in\mathrm{L}(D,Y)$, then, due to the properties of $f$ and $\eta$, the function $\widehat{\Pi}\colon{\mathbb{R}}_{+}\times\widetilde{M}\to W$ ( $W=Y_{1}\mathop{\dot{+}}Y_{2\Sigma}$}) is continuous and satisfies locally a Lipschitz condition with respect to $w$ on ${\mathbb{R}}_{+}\times\widetilde{M}$. By [29, p. 9, Theorem 1] we infer that for each initial point $(t_{0},x_{0})\in L_{0}$ such that $(P_{1}+P_{2\Sigma})x_{0}\in M_{1,2\Sigma}$ and $P_{20}x_{0}\in M_{20}$ the IVP for equation (4.14) with the initial condition
@@
has a unique solution $w(t)$ on some interval $J\subset{\mathbb{R}}_{+}$ that contains $t_{0}$. Using [29, p. 16-17, Corollary], we also infer that the solution $w(t)$ can be extended over a maximal interval of existence $t_{0},t_{max})$ ($t_{max}\leq\infty$}) of $w(t)$ in $\widetilde{M}$, and the extended solution is unique. Notice that the function $x(t)=\widetilde{A}^{-1}w(t)+\widehat{\eta}(t,w(t))$ is continuous and $Ax(t)=(Q_{1}+Q_{2\Sigma})w(t)=w(t)$ has the continuous derivative $\dfrac{dAx}{dt}(t)=\widehat{\Pi}(t,w(t))$ on $t_{0},t_{max})$ (for $t_{0}=0$, $Ax(t)$ has the derivative on the right at $t_{0}$). Thus, the IVP (1.1), (4.8) has the unique solution
@@
Note that from [19, Lemma 3.3] it follows that $\lim\limits_{t\to t_{max}-0}z(t)$ does not exist (the same follows from [29, p. 17, the proof of Corollary]), i.e., there is no point $z_{*}\in Z$ such that $\lim\limits_{t\to t_{max}-0}z(t)=z_{*}$.
@@
Let $(t_{0},x_{0})\in L_{0}$ is an arbitrary initial point. Since all the conditions of Lemma 4.3 (consequently, the conditions of Theorem 4.1) are satisfied, then the IVP (1.1), (4.8) has the unique solution $x(t)=\widetilde{A}^{-1}w(t)+\widehat{\eta}(t,w(t))$} on the maximal interval of existence $t_{0},t_{max})$ (see the proof of Theorem 4.1). Here $w(t)$ is the solution of the IVP (4.14), (4.15) on $t_{0},t_{max})$, $\widetilde{A}^{-1}\in\mathrm{L}(Y,D)$ and $\widehat{\eta}\in C({\mathbb{R}}_{+}\times W,D_{20})$ is the function defined in (4.14). Also, it follows from the proof of Lemma 4.3 that $\lim\limits_{t\to t_{max}-0}\|w(t)\|=\infty$ if $t_{max}<\infty$.
@@
The existence of the maximal interval of existence $t_{0},t_{max})$ of the solution $z(t)$ in $D$ follows from [19, Lemma 3.2]. From [19, Lemma 3.3] (or [29, p. 17, the proof of Corollary]) and the fact that the solution cannot leave $D$, we conclude that there is no point $z_{*}\in\overline{D}$ such that $\lim\limits_{t\to t_{max}-0}z(t)=z_{*}$. In addition, if $D$ is bounded, then due to [19, Theorem 2 and its Corollary] the solution exists on $t_{0},\infty)$. Also, note that if $\sup\limits_{t\,\in\,[t_{0},t_{max}),\,z\,\in\,D}\|H(t,z)\|<\infty$, then there exists $\lim\limits_{t\to t_{max}-0}z(t)\in\partial D$.

### v2

@@ 3 Nonlocal (global) implicit function theorems
The proof of the existence of a unique solution of equation (3.1) with respect to $y$ in a certain neighbourhood of a point $(t^{\prime},x^{\prime},y^{\prime})$ satisfying the equation is similar to the proof of the implicit function theorem without the requirement of differentiability [30, p. 420]. The Lipschitz continuity of the implicit function with respect to $x$ follows from the local Lipschitz continuity of $F$ with respect to $(x,y)$ and the contractivity of the mapping $\widehat{F}$ defined in (3.3) below. Note that $U_{r_{1}}(t^{\prime})$ can be a semi-open interval, as indicated in the introduction. In general, the idea of the proof is the same as for [11, Lemma 3.1]. However, since in [11] nonregular (singular) semilinear DAEs in finite-dimensional spaces were considered, then the form of components of the phase variable $x$, as well as direct decompositions of spaces and the associated projectors, is different from that of the present paper. The proof of [11, Lemma 3.1] is quite cumbersome due to the larger number of arguments of an implicit function and some features of the problem, so it will be easier to follow the proof of the lemma given above. Thus, for clarity, we provide the complete proof of the present lemma.
@@
It follows from the foregoing that, by the Banach’s fixed point theorem and the theorem on continuous dependence of a fixed point on a parameter (see, e.g., [27, Theorems 46, $46_{2}$]), the mapping $\widehat{F}(t,x,y)$ as a function of $y$, depending on the parameter $(t,x)$, has a unique fixed point $\upsilon_{(t,x)}=\upsilon(t,x)$ (i.e., $\widehat{F}(t,x,\upsilon(t,x))=\upsilon(t,x)$) in $\overline{U_{\tilde{\varepsilon}}}(y^{\prime})$ for each $(t,x)\in\overline{U_{\delta_{*}}}(t^{\prime},x^{\prime})$, which satisfies the equality $\upsilon(t^{\prime},x^{\prime})=y^{\prime}$ and depends continuously on $(t,x)$. The continuity of the function $\upsilon\colon\overline{U_{\delta_{*}}}(t^{\prime},x^{\prime})\to\overline{U_{\tilde{\varepsilon}}}(y^{\prime})$ is proved in the same way as in [27, Theorem $46_{2}$].
@@
Choose an arbitrary fixed point $(t^{\prime},x^{\prime})\in{\mathcal{T}}\times M_{X}$. Then, due to conditions 1 and 2, there exists a unique $y^{\prime}\in M_{Y}$} such that $(t^{\prime},x^{\prime},y^{\prime})$ satisfies (3.1) and the operator $\partial_{y}F(t^{\prime},x^{\prime},y^{\prime})$ has the bounded inverse $\partial_{y}F(t^{\prime},x^{\prime},y^{\prime})]^{-1}\in\mathrm{L}(Z,Y)$. Using the theorems on the existence and differentiability of an implicit function [27, Theorems 25, 28], we obtain that there exist open neighborhoods $U_{\delta_{1}}(t^{\prime})$, $U_{\delta_{2}}(x^{\prime})$ and $U_{\varepsilon}(y^{\prime})$ and a unique function $\upsilon_{t^{\prime},x^{\prime}}\in C\big(U_{\delta_{1}}(t^{\prime})\times U_{\delta_{2}}(x^{\prime}),U_{\varepsilon}(y^{\prime})\big)$ such that $\upsilon_{t^{\prime},x^{\prime}}(t^{\prime},x^{\prime})=y^{\prime}$, $\upsilon_{t^{\prime},x^{\prime}}(t,x)$ is a solution of equation (3.1) with respect to $y$, and it is continuously differentiable with respect to $x$ on $U_{\delta_{1}}(t^{\prime})\times U_{\delta_{2}}(x^{\prime})$. Recall that $U_{\delta_{1}}(t^{\prime})$ can be a semi-open interval in $\mathcal{T}$}. Generally, the implicit function theorems [27] assume that the set of variables is open, and therefore, to prove the existence of an implicit function when the interval $U_{\delta_{1}}(t^{\prime})$ is semi-open, the fixed point theorems [27, Theorems 46, $46_{2}$] as well as the proofs of the implicit function theorems [27, Theorems 25, 28] (which in turn use the same fixed point theorems) are used.
@@
Let us show this. Since $F(t,x,y)$ has a continuous partial derivatives with respect to $(x,y)$ on ${\mathcal{T}}\times D_{X}\times D_{Y}$, then it is locally Lipschitz continuous with respect to $(x,y)$. Choose an arbitrary fixed point $(t^{\prime},x^{\prime},y^{\prime})\in{\mathcal{T}}\times M_{X}\times M_{Y}$ at which equation (3.1) holds, and take the operator $W_{t^{\prime},x^{\prime},y^{\prime}}=\partial_{y}F(t^{\prime},x^{\prime},y^{\prime})$ as the operator $W_{t^{\prime},x^{\prime},y^{\prime}}$ appearing in condition 2 of Theorem 3.1. Then, due to condition 2 of Theorem 3.2, there exist open neighborhoods $U_{\delta}(t^{\prime},x^{\prime})\subset{\mathcal{T}}\times D_{X}$ and $U_{\varepsilon}(y^{\prime})\subset D_{Y}$ such that for each $(t,x)\in\overline{U_{\delta}(t^{\prime},x^{\prime})}$, $y^{1},\,y^{2}\in\overline{U_{\varepsilon}(y^{\prime})}$ the following holds: $\big\|F(t,x,y^{1})-F(t,x,y^{2})-W_{t^{\prime},x^{\prime},y^{\prime}}\big[y^{1}-y^{2}\big]\big\|\leq\int\limits_{0}^{1}\big\|\partial_{y}F\big(t,x,y^{2}+\theta(y^{1}-y^{2})\big)-W_{t^{\prime},x^{\prime},y^{\prime}}\big\|d\theta\,\|y^{1}-y^{2}\|\leq k(\delta,\varepsilon)\|y^{1}-y^{2}\|$, where $k(\delta,\varepsilon)=\sup\limits_{(t,x)\in\overline{U_{\delta}(t^{\prime},x^{\prime})},\,y^{\prime\prime}\in\overline{U_{\varepsilon}(y^{\prime})}}\big\|\partial_{y}F(t,x,y^{\prime\prime})-W_{t^{\prime},x^{\prime},y^{\prime}}\big\|\to 0$ as $\delta,\varepsilon\to 0$ since $\partial_{y}F$ is continuous at $(t^{\prime},x^{\prime},y^{\prime})$ (cf. [30, p. 422, the proof of the theorem]). Hence, taking into account that $W_{t^{\prime},x^{\prime},y^{\prime}}$ has a bounded inverse, condition 2 of Theorem 3.1 holds.
@@
The proof is similar to the proof of Theorem 3.2, except that the theorem on higher derivatives of an implicit function [27, Theorem 31] is also used here. ∎
Another nonlocal (global) implicit function theorems have been presented, e.g., in [1, 22, 4] (see also references therein). In [1], a nonlocal implicit function theorem for a closed mapping with a parameter from one Asplund space to another has been obtained in terms of the regular coderivative of the mapping. In [22], nonlocal implicit function theorems for mappings in Banach spaces has been obtained by using the concept of “continuation property” and local mapping relations. For mappings in real finite-dimensional spaces, a nonlocal implicit function theorem is presented in [4] with the use of the maximal rank condition.
@@ 4.1 First approach
In this subsection, we consider the case when the component $x_{20}=P_{20}x$ can be defined as an implicit function from equation (4.2) given below (i.e., when the projection $Q_{2*}\big[f(t,x)-B(x_{2\Sigma}+x_{20})\big]$ depends on $x_{20}$; for clarity, notice that $Q_{2*}Bx_{20}=Bx_{20}^{(1)}$ where $x_{20}^{(1)}=P_{20}^{(1)}x$). In the case when the pencil $P(\lambda)$ has index 1, this always holds, since $D_{20}=D_{2}$, $x_{20}=x_{2}=P_{2}x$, $x_{2\Sigma}=0$, $Q_{2*}=Q_{2}$ ($P_{2\Sigma}=0$, $Q_{2\Sigma}=0$) and $B_{2}=(Q_{2}B)\big|_{D_{2}}$ has the inverse $B_{2}^{-1}\in\mathrm{L}(Y_{2},D_{2})$.
@@
Remark 4.1.
In what follows, we assume that $D_{B}\supseteq D_{A}$ (then $D=D_{A}$), but this condition is not necessary (we make this assumption to simplify the presentation of further results). If we do not require that $D_{B}\supseteq D_{A}$, then it is necessary to require that $\overline{D}=X$ and to replace $A$ by the restricted operator $A\big|_{D}\colon D\to Y$ and, accordingly, the semi-inverse operator $A^{(-1)}$ by the semi-inverse $\big(A\big|_{D}\big)^{(-1)}$ below (a semi-inverse operator [6] is called an algebraic generalized inverse and, in Banach spaces, a topological generalized inverse in [20]). Then the results of Section 4 remain valid.
@@
$F_{2*}(t,x_{1},x_{2\Sigma},x_{20})=0,\quad F_{2*}(t,x_{1},x_{2\Sigma},x_{20}):=Q_{2*}f(t,x)-Q_{2*}B(x_{2\Sigma}+x_{20}),$ (4.4)
where the operator $\widetilde{A}^{-1}\in\mathrm{L}(Y,X)$ is the inverse of
@@
and $A^{(-1)}\in\mathrm{L}(Y,X)$ is the semi-inverse operator to $A$, which can be obtained by $A^{(-1)}=\widetilde{A}^{-1}(Q_{1}+Q_{2\Sigma})$} (or by using the relations mentioned below). The existence and continuity of $\widetilde{A}^{-1}$ is proved in a similar way as the Schmidt lemma (see [30, p. 232] or [31]). It is readily verified that the operator $\widetilde{A}^{-1}=A^{(-1)}+\sum\limits_{i=1}^{n}\langle\cdot,q_{i}^{m_{i}}\rangle\varphi_{i}^{1}$ is the inverse of $\widetilde{A}$. Notice that $Y_{1}=\widetilde{A}D_{1}=AD_{1}$,  $Y_{2\Sigma}=\widetilde{A}D_{2\Sigma}=AD_{2\Sigma}$,  $\widetilde{A}^{-1}Ax=(P_{1}+P_{2\Sigma})x$, $x\in D$, and $A\widetilde{A}^{-1}y=(Q_{1}+Q_{2\Sigma})y$, $y\in Y$. In addition, $Q_{1}\widetilde{A}x=Q_{1}Ax=AP_{1}x=\widetilde{A}P_{1}x$ and $Q_{2\Sigma}\widetilde{A}x=Q_{2\Sigma}Ax=AP_{2\Sigma}x=\widetilde{A}P_{2\Sigma}x$, $x\in D$.
The semi-inverse operator $A^{(-1)}$ is defined by the relations
$\begin{split}&A^{(-1)}A=(P_{1}+P_{2\Sigma})\big|_{D_{A}},\quad AA^{(-1)}=Q_{1}+Q_{2\Sigma},\\ &A^{(-1)}=(P_{1}+P_{2\Sigma})A^{(-1)}\quad(\text{or }A^{(-1)}=A^{(-1)}(Q_{1}+Q_{2\Sigma})\,).\end{split}$ (4.6)
Note that from these relations the equalities $AA^{(-1)}A=A$ and $A^{(-1)}AA^{(-1)}=A^{(-1)}$ follows, and the converse is also true with respect to the direct decompositions (2.10), (2.11) and the corresponding projectors.
@@
where $\widetilde{A}=A+BP_{2}=Q_{1}A+Q_{2}B$ coincide with the operator $G=Q_{1}A+Q_{2}B$ from [32, Subsection 3.3] which has the inverse $G^{-1}=\widetilde{A}^{-1}\in\mathrm{L}(Y,X)$.
@@
The “algebraic part” (4.2) of the ADAE generates the manifold
@@
Since $\widetilde{A}^{-1}\in\mathrm{L}(Y,X)$, $Q_{1}+Q_{2\Sigma}\in\mathrm{L}(Y)$, $(Q_{1}+Q_{2\Sigma})B\widetilde{A}^{-1}\in\mathrm{L}(Y)$ and $(Q_{1}+Q_{2\Sigma})BP_{20}\in\mathrm{L}(D,Y)$, then, due to the properties of $f$ and $\eta$, the function $\widehat{\Pi}\colon{\mathbb{R}}_{+}\times\widetilde{M}\to W$ ( $W=Y_{1}\mathop{\dot{+}}Y_{2\Sigma}$}) is continuous and satisfies locally a Lipschitz condition with respect to $w$ on ${\mathbb{R}}_{+}\times\widetilde{M}$. By [28, p. 9, Theorem 1] we infer that for each initial point $(t_{0},x_{0})\in L_{0}$ such that $(P_{1}+P_{2\Sigma})x_{0}\in M_{1,2\Sigma}$ and $P_{20}x_{0}\in M_{20}$ the IVP for equation (4.15) with the initial condition
@@
has a unique solution $w(t)$ on some interval $J\subset{\mathbb{R}}_{+}$ that contains $t_{0}$. Using [28, p. 16-17, Corollary], we also infer that the solution $w(t)$ can be extended over a maximal interval of existence $t_{0},t_{max})$ ($t_{max}\leq\infty$}) of $w(t)$ in $\widetilde{M}$, and the extended solution is unique. Notice that the function $x(t)=\widetilde{A}^{-1}w(t)+\widehat{\eta}(t,w(t))$ is continuous and $Ax(t)=(Q_{1}+Q_{2\Sigma})w(t)=w(t)$ has the continuous derivative $\dfrac{dAx}{dt}(t)=\widehat{\Pi}(t,w(t))$ on $t_{0},t_{max})$ (for $t_{0}=0$, $Ax(t)$ has the derivative on the right at $t_{0}$). Thus, the IVP (1.1), (4.9) has the unique solution
@@
Note that from [19, Lemma 3.3] it follows that $\lim\limits_{t\to t_{max}-0}z(t)$ does not exist (the same follows from [28, p. 17, the proof of Corollary]), i.e., there is no point $z_{*}\in Z$ such that $\lim\limits_{t\to t_{max}-0}z(t)=z_{*}$.
@@
Let $(t_{0},x_{0})\in L_{0}$ is an arbitrary initial point. Since all the conditions of Lemma 4.3 (consequently, the conditions of Theorem 4.1) are satisfied, then the IVP (1.1), (4.9) has the unique solution $x(t)=\widetilde{A}^{-1}w(t)+\widehat{\eta}(t,w(t))$} on the maximal interval of existence $t_{0},t_{max})$ (see the proof of Theorem 4.1). Here $w(t)$ is the solution of the IVP (4.15), (4.16) on $t_{0},t_{max})$, $\widetilde{A}^{-1}\in\mathrm{L}(Y,X)$ and $\widehat{\eta}\in C({\mathbb{R}}_{+}\times W,D_{20})$ is the function defined in (4.15). Also, it follows from the proof of Lemma 4.3 that $\lim\limits_{t\to t_{max}-0}\|w(t)\|=\infty$ if $t_{max}<\infty$.
@@
The existence of the maximal interval of existence $t_{0},t_{max})$ of the solution $z(t)$ in $D$ follows from [19, Lemma 3.2]. From [19, Lemma 3.3] (or [28, p. 17, the proof of Corollary]) and the fact that the solution cannot leave $D$, we conclude that there is no point $z_{*}\in\overline{D}$ such that $\lim\limits_{t\to t_{max}-0}z(t)=z_{*}$. In addition, if $D$ is bounded, then due to [19, Theorem 2 and its Corollary] the solution exists on $t_{0},\infty)$. Also, note that if $\sup\limits_{t\,\in\,[t_{0},t_{max}),\,z\,\in\,D}\|H(t,z)\|<\infty$, then there exists $\lim\limits_{t\to t_{max}-0}z(t)\in\partial D$.
@@
Remark 4.7.
In the above theorems, lemmas and corollaries one can use the semi-inverse operator $A^{(-1)}$ defined by (4.6) instead of the inverse $\widetilde{A}^{-1}$ of (4.5), taking into account that $A^{(-1)}=\widetilde{A}^{-1}(Q_{1}+Q_{2\Sigma})$},  $\widetilde{A}^{-1}\widetilde{M}=A^{(-1)}\widetilde{M}$  ($\widetilde{M}\subseteq Y_{1}\mathop{\dot{+}}Y_{2\Sigma}$),  $\widetilde{A}(x_{1}+x_{2\Sigma})=A(x_{1}+x_{2\Sigma})$,  $\widetilde{A}^{-1}w=A^{(-1)}w$  ($w\in Y_{1}\mathop{\dot{+}}Y_{2\Sigma}$),  $\widetilde{A}\,\Pi(t,x)=A\,\Pi(t,x)$,  $\widetilde{A}^{-1}\widetilde{\Omega}=A^{(-1)}\widetilde{\Omega}$  ($\widetilde{\Omega}\subseteq Y_{1}\mathop{\dot{+}}Y_{2\Sigma}$).